### v1

# Limits to extreme event forecasting in chaotic systems

Abstract
Predicting extreme events in chaotic systems, characterized by rare but intensely fluctuating properties, is of great importance due to their impact on the performance and reliability of a wide range of systems. Some examples include weather forecasting, traffic management, power grid operations, and financial market analysis, to name a few. Methods of increasing sophistication have been developed to forecast events in these systems. However, the boundaries that define the maximum accuracy of forecasting tools are still largely unexplored from a theoretical standpoint. Here, we address the question: What is the minimum possible error in the prediction of extreme events in complex, chaotic systems? We derive lower bounds for the minimum probability of error in extreme event forecasting using the information-theoretic Fano’s inequality. The limits obtained are universal, in that they hold regardless of the modeling approach: from traditional linear regressions to sophisticated neural network models. The approach also allows us to assess whether reduced-order models are operating near their theoretical maximum performance or if further improvements are theoretically possible.
keywords:
[…]
1 Introduction
Extreme events, characterized by rare but intensely fluctuating properties, are ubiquitous in both engineering system and natural phenomena Albeverio2006 . For instance, turbulent gusts over an aircraft can result in bumpy flights Etkin1981 , severe weather can disrupt communication systems Rahmstorf2011 , rare but large cascades in electrical power grids may lead to failures Schaefer2018 , and sudden increases in traffic flow can trigger network paralysis Kumar2020 . In these scenarios, the real-time prediction of extreme events is crucial for enabling proactive measures to avert potential issues Farazmand2019 ; Sapsis2021 . By accurately forecasting the extreme states of dynamical systems, we can mitigate adverse effects, reduce downtime, and prevent failures. In this study, we present the lower bound on the accuracy of extreme event prediction using the framework of information theory. The bound obtained is a fundamental limit –independent of the modeling approach– that arises from the finite amount of information the observed state contains about the extreme event.
A variety of methods have been employed to predict extreme events in time series of chaotic dynamical systems. Some of the approaches that have proven effective include nonlinear dynamics estimation based on the Koopman operator theory Otto2021 and Takens embedding theorem Asch2022 , along with machine learning techniques, such as support vector machines Mukherjee1997 , singular spectrum analysis and the maximum entropy method Ghil2002 . Advanced deep learning methods, including auto-encoders Racah2017 , long short-term memory networks Vlachas2018 , and reservoir computing Pammi2023 have also been instrumental to devise forecasting models for chaotic systems with high-dimensional attractors.
Despite the significant advancements described above, the inherent nature of chaos continues to impose limits on the accuracy of models for extreme event forecasting. The error of prediction in chaotic dynamical systems stem from three primary sources Smith2001 . First, the model might not accurately represent the physical reality. Second, the observable variables may not capture all the relevant degrees of freedom present in the dynamical system. Third, the initial conditions required for forecasting might not be precisely known.
Improvements in the prediction of extreme events can be achieved either by enhancing models to better represent the physics, gaining access to more observables, or reducing uncertainty in the initial conditions. Eliminating modeling errors is theoretically possible, given the knowledge of a set of governing equations that reflect the underlying dynamics of the observed system. However, accessing variables beyond what is currently observable may be limited by experimental/computational constraints. Additionally, no feasible approach can completely eliminate prediction errors caused by uncertainty in the initial conditions. Even with highly precise measurements, minor errors in the initial state eventually amplify, compromising the accuracy of the forecast for long times (Lorenz1963, ). Here, our focus is not on developing superior models for extreme event prediction. Instead, we pose the fundamental question: what is the theoretically maximum achievable accuracy in extreme event prediction regardless of the modeling approach and source of error?
2 Methods
Consider the quantity of interest $Q(t)$ from a chaotic system, where $t$ is time. The extreme event indicator $E(t)$ is defined as
$E(t)=\begin{cases}1&\text{ if }Q(t)\geqslant\eta,\\ 0&\text{otherwise},\end{cases}$ (1)
where $\eta$ is the threshold for extreme event detection. The vector of observable variables is $\boldsymbol{\hat{Q}}(t)=[\hat{Q}_{1}(t),\hat{Q}_{2}(t),...]$, which contains the accessible information about the system. One could forecast the extreme event $E(t)$ at time $t$ using the finitely precise observable in the past, $\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{Q}}(t-\delta t)\pm\delta\boldsymbol{\hat{Q}}$, as the input to the model $\hat{f}$ such that $\hat{E}(t)=\hat{f}\left(\boldsymbol{\hat{Q}}^{-}\right)$, where $\hat{E}$ is the modeled extreme event indicator (which might differ from $E$), $\delta t>0$ is the time horizon for extreme event forecasting, and $\delta\boldsymbol{\hat{Q}}$ is the uncertainty in the state $\boldsymbol{\hat{Q}}$. The performance of the model can be evaluated using the probability of error $P_{e}(\boldsymbol{\hat{Q}}^{-},\hat{f})=\text{Probability}(\hat{E}\neq E)=P(\hat{f}(\boldsymbol{\hat{Q}}^{-})\neq E)$.
Mispredicted extreme events can manifest as either false positives, $P(\hat{E}=1,E=0)$, or false negatives, $P(\hat{E}=0,E=1)$. However, these two types of errors can bear significantly different consequences. For instance, incorrectly predicting a hurricane (false positive) might be inconvenient, but failing to predict one (false negative) can be catastrophic. To accurately reflect the distinct impact of false positive and negative, we introduce the cost-sensitive probability of error:
$P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})=c^{+}P(\hat{E}=1,E=0)+c^{-}P(\hat{E}=0,E=1),$ (2)
where $c^{+}$ and $c^{-}$ are the false positive and negative weighting factors, respectively.
The goal is to estimate the minimum cost-sensitive probability of error given the observable state $\boldsymbol{\hat{Q}}^{-}$ over all possible models $\hat{f}$,
$P_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-})=\min_{\hat{f}}P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f}).$ (3)
The factors $c^{-}$ and $c^{+}$ are scaled such that $21/c^{+}+1/c^{-}=$ 2 to ensure that the range of $P^{c}_{e,\min}(\boldsymbol{\hat{Q}}^{-})$ is [0, 0.5]. The key idea is that the prediction of extreme events can be intuitively understood as an information transmission process, where information from the current observable state is conveyed to predict the future state (LozanoDuran2022, ). If the forecast is treated as a noisy channel, then the Fano’s inequality (Fano1961, ) provides the foundations for deriving a bound on the minimum probability of error in the transmission of discrete messages. First, we measure the uncertainty in the extreme event indicator $E$ after knowing the present observable $\boldsymbol{\hat{Q}}^{-}$ using the cost-sensitive, conditional Shannon entropy (Shannon1948, ; Zhao2013, )
$H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ $=\sum_{\boldsymbol{\hat{q}}^{-}}h^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),$ (4)
where $\boldsymbol{\hat{q}}^{-}$ is a particular state for $\boldsymbol{\hat{Q}}^{-}$, $P(E\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ is the conditional probability of $E$ on $\boldsymbol{\hat{Q}}^{-}$, and $h^{c}(p)=h\left(\min\{c^{-}p,c^{+}(1-p)\}\right)$ is the cost-sensitive binary entropy function, where $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$. For $c^{+}=c^{-}=1$, $H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ is equal to the classic Shannon conditional entropy. Equation (4) quantifies the additional information required to determine the outcome of $E$ given the information in $\boldsymbol{\hat{Q}}^{-}$ accounting for the weighting factors $c^{-}$ and $c^{+}$. With that, the minimum cost-sensitive probability of error attainable by any model is [see Appendix A]
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=h^{-1}\left(H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right).$ (5)
The lower bound given by Eq. (5) is the consequence of the unavoidable uncertainty intrinsic to chaotic systems. This uncertainty arises from unobserved variables and errors in the initial condition values, which transcend the predictive capabilities of any model. It is worth noting that while the bound in Eq. (5) is utilized here to establish limits for predicting extreme events, it is generally applicable to the prediction of any binary events.
The minimum probability of error in the extreme event forecaster can be decomposed into three sources, $P^{c}_{e}(\boldsymbol{\hat{Q}}^{-},\hat{f})=P^{c}_{e,I}+P^{c}_{e,O}+P^{c}_{e,M}$ given by
$P^{c}_{e,I}$ $=P_{e,\min}^{c}(\boldsymbol{Q}^{-}),$ (6)
$P^{c}_{e,O}$ $=P_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-})-P_{e,\min}^{c}(\boldsymbol{Q}^{-}),$
$P^{c}_{e,M}$ $=P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})-P_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-}),$
where $\boldsymbol{Q}^{-}=\boldsymbol{Q}(t-\delta t)\pm\delta\boldsymbol{Q}$ contains all the degrees of freedom governing the system (i.e., absolute observability) with finite precision, $P^{c}_{e,I}$ represents the probability of error solely due to uncertainty in the initial conditions, $P^{c}_{e,O}$ denotes the probability of error caused by missing information from unobserved variables, and $P^{c}_{e,M}$ is the probability of error attributable to a suboptimal model.
In the following, we demonstrate the application of our results in two distinct scenarios: the Rössler system and the Kolmogorov flow. The Rössler system offers a simple case for studying extreme events in a chaotic system where all variables can be observed. We use this case to illustrate the classification of errors from Eq. (6). On the other hand, the Kolmogorov flow, characterized by complex, multi-scale interactions among numerous degrees of freedom, represents the dynamics of extreme events found in realistic systems. This case is used to demonstrate the effect of risk-sensitive analysis.
3 Applications
3.1 Rössler system
The Rössler system with state variables $\boldsymbol{Q}=[\theta_{1},\theta_{2},\theta_{3}]$ is governed by the ordinary differential equation:
$\frac{\mathrm{d}\theta_{1}}{\mathrm{d}t}$ $=-\theta_{2}-\theta_{3},$ (7)
$\frac{\mathrm{d}\theta_{2}}{\mathrm{d}t}$ $=\theta_{1}+a\theta_{2},$
$\frac{\mathrm{d}\theta_{3}}{\mathrm{d}t}$ $=b+\theta_{3}(\theta_{1}-c),$
with parameters $a=0.1,b=0.1$, and $c=14$. The extreme events are based on $\theta_{3}$, which exhibits rare excursions of intense magnitude. Figure 1(a) shows the trajectory of the Rössler system in the three-dimensional phase space. The extreme event indicator is defined as
$E(t)=\begin{cases}1&\text{ if }\theta_{3}(t)\geqslant\bar{\theta}_{3}+3\sigma_{\theta_{3}},\\ 0&\text{ otherwise, }\end{cases}$ (8)
where $\bar{\theta}_{3}$ and $\sigma_{\theta_{3}}$ are the mean and standard deviation of ${\theta_{3}}$ over time. Figure 1(b) contains a fragment of the time history of $\boldsymbol{Q}$ and the extreme event indicator $E$. We investigate the case with balanced risk $c^{+}=c^{-}=1$.
We define the normalized probability of error as $\bar{P}_{e}^{c}=P_{e}^{c}/P(E=1)$, where $P(E=1)$ is the probability of extreme events such that $\bar{P}_{e,\min}^{c}\rightarrow 1$ for $\delta t\rightarrow\infty$. Figure 2 shows the normalized minimum probability of error as a function of $\delta t$. Two scenarios are considered. In the first case (represented by a solid line), we assume the only observable variable is $\hat{Q}^{-}=\theta_{3}(t-\delta t)+\delta\theta_{3}$, i.e., $P_{e,\min}^{c}(\hat{Q}^{-})$, where the uncertainty $\delta\theta_{3}$ is 10% of $\sigma_{\theta_{3}}$. In the second scenario (dashed line), it is presumed that knowledge of the full state $\boldsymbol{Q}^{-}$ is available, i.e., $P_{e,\min}^{c}(\boldsymbol{Q}^{-})$. The region beneath each curve corresponds to models that are unattainable, while the region above represents models that are underperforming. The two curves also allow us to discern between errors from uncertainty in initial conditions ($P^{c}_{e,I}$) and errors from the unobserved variables $\theta_{1}$ and $\theta_{2}$ ($P^{c}_{e,O}$). Over time, both cases converge to $\bar{P}_{e}^{c}\rightarrow 1$ given the chaotic nature of the system. This convergence is slower for $P_{e,\min}^{c}(\boldsymbol{Q}^{-})$, as errors are solely due to uncertainties in the initial condition.
Figure 1: (a) Trajectory of the Rössler system. (b) Extraction of time series of $\theta_{1}$, $\theta_{2}$, $\theta_{3}$ and $D$ in the Rössler system. Although not shown, the whole time-span of the signals is 10,000 time units.
Figure 2: Normalized probability of error for extreme events prediction in the Rössler system. $P^{c}_{e,I}$: error due to uncertainty in the initial conditions; $P^{c}_{e,O}$: error caused by unobserved variables; $P^{c}_{e,M}$: error attributable to a suboptimal model.
We also consider a logistic regression model that uses $\theta_{3}$ as input and is trained for different forecast times $\delta t$ (also included in Fig. 2). The probability of error for the logistic regression model, $P_{e}^{c}(\hat{Q}^{-},\hat{f}^{LR})$, enables the quantification of the model error ($P^{c}_{e,M}$) for this specific case. There is a significant gap between $P^{c}_{e,M}$ and the error bounds, implying that potential model improvements are possible. As $\delta t$ increases, $P^{c}_{e,M}$ rapidly becomes the predominant source of error: $P^{c}_{e,I}$ is minor compared to $P^{c}_{e,O}$ and $P^{c}_{e,M}$, indicating that the missing variables and the suboptimal model have a more significant impact on the accuracy of the forecast compared to errors in the initial condition.
3.2 Kolmogorov flow
Next, we evaluate the cost-sensitive error bounds for forecasting intense energy dissipation events in a turbulent flow Farazmand2017 . The case considered is the Kolmogorov flow: a high-dimensional, chaotic dynamical system described by the two-dimensional Navier-Stokes equations and driven by monochromatic body forcing Arnold1960 :
$\frac{\partial u_{i}}{\partial t}$ $=-\frac{\partial(u_{i}u_{j})}{\partial x_{j}}-\frac{\partial p}{\partial x_{i}}+\frac{1}{\text{Re}}\frac{\partial^{2}u_{i}}{\partial x_{k}\partial x_{k}}+f_{i},\quad i=1,2,$ (9)
$\frac{\partial u_{i}}{\partial x_{i}}$ $=0,$ (10)
where repeated indices imply summation, $u_{i}(x_{1},x_{2},t)$ is the velocity $i$-th velocity component, $p$ is the pressure, $x_{1}$ and $x_{2}$ are the spatial coordinates, and $f_{i}$ is the forcing with $f_{1}=\sin(4x_{2})$ and $f_{2}=0$. The flow setup is characterized by the Reynolds number $\text{Re}\approx 100$ for which the flow exhibits intermittent bursts of dissipation events Zeff2003 . For our analysis, we use data from Farazmand and Sapsis Farazmand2017 , obtained by numerically resolving all the scales of the problem in a doubly periodic box with side $2\pi$ and $256^{2}$ spatial Fourier modes. Figure 3(a) shows the velocity amplitude $|\boldsymbol{u}(x_{1},x_{2},t)|$ at a given time.
[…]
$E(t)=\begin{cases}1&\text{ if }D(t)\geqslant\bar{D}+1.5\sigma_{D},\\ 0&\text{ otherwise},\end{cases}$ (11)
where $\bar{D}$ and $\sigma_{D}$ are the mean and standard deviation of ${D}$ over time. The observable chosen is the magnitude of the spatial Fourier mode of velocity vector $\boldsymbol{u}=[u_{1},u_{2}]$ corresponding to the wavenumber $[1,0]$, which is denoted by $\hat{Q}=\left|\boldsymbol{\breve{u}}_{1,0}\right|$. An excerpt of $\hat{Q}$ and $D$ extracted from the full time history is presented in Fig. 3(b). The time is non-dimensionalized by $t_{e}=(\text{Re}\bar{D})^{-1/2}$.
Figure 3: (a) The instantaneous velocity amplitude $|\boldsymbol{u}|$ for the Kolmogorov flow. (b) Time history of $D(t)$ and $|\boldsymbol{\breve{u}}_{1,0}|(t)$.
Three cases are investigated: $c^{-}$ = 1, 1.5, and 2. The first case ($c^{-}$ = 1) penalizes false positives and false negatives equally. The other two cases assign a higher penalty to false negatives, such that the cost of failing to predict an extreme event is two to three times the cost of incorrectly predicting a non-extreme event. We denote the minimum probability of error for each case as $P_{e,\min}^{c^{-}=1}(\hat{Q}^{-})$, $P_{e,\min}^{c^{-}=1.5}(\hat{Q}^{-})$, and $P_{e,\min}^{c^{-}=2}(\hat{Q}^{-})$, respectively. Figure 4 shows the results as a function of $\delta t$. The minimum cost-sensitive probability of error is normalized as $\bar{P}_{e,\min}^{c}=P_{e,\min}^{c}/\min\{\left(c^{-}\cdot P(E=1),c^{+}\cdot(1-P(E=1)\right)\}$ such that $\bar{P}_{e,\min}^{c}\rightarrow 1$ for $\delta t\rightarrow\infty$.
Figure 4: Normalized, cost-sensitive minimum probability of error for extreme event prediction in the Kolmogorov flow.
The results from Fig. 4 show that increasing $c^{-}$ reduces the normalized minimum probability of error, making the predictions less challenging. This trend is particular to the Komogorov flow and the chosen variables, and other systems can exhibit different behavior. Figure 4 also illustrate two interesting characteristics of $\bar{P}_{e,\min}^{c}(\hat{Q}^{-})$ for the three $c^{-}$ values considered. First, $\bar{P}_{e,\min}^{c}(\hat{Q}^{-})$ is not initially zero when $\delta t=0$. This situation arises because $D$ was not considered an observed variable, leading to uncertainty in $E$ even at $\delta t=0$. The second interesting observation is that $\bar{P}_{e,\min}^{c}$ reaches its lowest value at $\delta t_{\min}=2.8t_{e}$. This observation can be understood by noting that in this system, energy is transferred among different scales due to nonlinear interactions until it is ultimately dissipated. This process occurs on a timescale comparable to $\delta t_{\min}$ (Farazmand2017, ), which explains the effectiveness of $\left|\boldsymbol{\breve{u}}_{1,0}\right|$ in predicting extreme dissipation events at that time lag. For times beyond this point, $\left|\boldsymbol{\breve{u}}_{1,0}\right|$ becomes less useful due to the chaoticity of turbulence.
4 Conclusions
In this study, we have used the information-theoretic Fano’s inequality to derive the lower bound for predicting extreme events in chaotic time series. The applications discussed here demonstrate the utility of the lower bound obtained (Eq. 5) as a tool for investigating the intrinsic limitations to forecast extreme events in chaotic systems. We have shown that this bound maintains its validity irrespective of the chosen modeling method and plays a crucial role in assessing whether models are functioning near their theoretical optimum. Furthermore, the bound is applicable to scenarios involving both balanced and unbalanced risks and can be used to investigate sources of error from initial conditions, unobserved variables, and suboptimal modeling assumptions (Eq. 6).
The use of Eq. (5) extends beyond the extreme events prediction presented here; it is broadly applicable to any binary classification of events, regardless of whether they are categorized as extreme/non-extreme. Equation (5) also lays the foundations for future extensions to forecasting continuous time signals by employing generalized versions of Fano’s inequality.
5 Declaration of competing interest
There is no conflict of interest between the authors and other persons organizations.
6 Data availability
All the data from this research are available upon request.
7 Acknowledgments
This work was supported by the National Science Foundation under Grant No. 2140775 and MISTI Global Seed Funds.
Appendix A Proof of information-theoretic bound
The minimum cost-sensitive probability of error achievable by any model based on the cost-sensitive uncertainty in $E$ conditioned to the observable $\boldsymbol{\hat{Q}}^{-}$ is
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=h^{-1}\left(H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)\leq P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f}).$ (12)
Proof.
The overall cost-sensitive probability of error is a weighted sum of the probability of error for each specific state of $\boldsymbol{\hat{Q}}^{-}$,
$P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})$ $=c^{-}P(\hat{E}=0,E=1)+c^{+}P(\hat{E}=1,E=0)$ (13)
$=\sum_{\boldsymbol{\hat{q}}^{-}}\biggl{(}c^{-}P(\hat{E}=0,E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$
$\qquad+c^{+}P(\hat{E}=1,E=0\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\biggr{)}P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),$
where we assume that $21/c^{-}+1/c^{+}=$ 2 to ensure that
$P_{e}^{e,\min}(\boldsymbol{\hat{Q}}^{-})\leq\min\left\{c^{-}P(E=1),c^{+}\left(1-P(E=1)\right)\right\}\leq\frac{c^{-}c^{+}}{c^{+}+c^{-}}=\frac{1}{2}.$ (14)
For each specific state of $\boldsymbol{\hat{Q}}^{-}$, the minimum probability of error is determined by
$c^{-}P(\hat{E}=0,E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})+c^{+}P(\hat{E}=1,E=0\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ (15)
=$=\begin{cases}c^{-}P(\hat{E}=0,E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),&\text{if }\hat{E}(\boldsymbol{\hat{q}}^{-})=0\\ c^{+}P(\hat{E}=1,E=0\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),&\text{if }\hat{E}(\boldsymbol{\hat{q}}^{-})=1\end{cases}$
=$=\begin{cases}c^{-}P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})-c^{-}\underbrace{P(\hat{E}=1,E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})}_{0},&\text{if }\hat{E}(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=0\\ c^{+}P(E=0\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})-c^{+}\underbrace{P(\hat{E}=0,E=0\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})}_{0},&\text{if }\hat{E}(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=1\end{cases}$
$\geq\min\left\{c^{-}P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),c^{+}P(E=0\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right\}$
$=\min\left\{c^{-}P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),c^{+}\left(1-P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)\right\}.$
Let us define the minimum probability of error at each state as
$I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=\min\left\{c^{-}P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),c^{+}\left(1-P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)\right\},$ (16)
applying Eq. (13) and Eq. (15) we have,
$P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})\geq\sum_{\boldsymbol{\hat{q}}^{-}}I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})].$ (17)
Given the concave function $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$ for $p\in[0,0.5]$, the Jensen’s inequality for the random variable $I(\boldsymbol{\hat{Q}}^{-})$ results in
$h\left(\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]\right)\geq\mathbb{E}\left[h\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right].$ (18)
On the left hand side of Eq. (18), noticing that $h(p)$ is a non-decreasing function in the range $[0,0.5]$ and utilizing Eq. (17),
$h\left(\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]\right)\leq h\left(P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})\right).$ (19)
The right hand side of Eq. (18) is
$\mathbb{E}\left[h\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right]=\sum_{\boldsymbol{\hat{q}}^{-}}h\left(I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ (20)
$=\sum_{\boldsymbol{\hat{q}}^{-}}h\left(\min\left\{c^{-}P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),c^{+}(1-P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}))\right\}\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$
$=\sum_{\boldsymbol{\hat{q}}^{-}}h^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=H^{c}(E\mid\boldsymbol{\hat{Q}}^{-}),$
where $H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ is defined as the cost-sensitive conditional entropy, and the cost-sensitive binary entropy function is given by
$h^{c}(p)=\begin{cases}h\left(c^{-}p\right)&\text{ for }p\in\left[0,\frac{1}{2c^{-}}\right),\\ h\left(c^{+}(1-p)\right)&\text{ for }p\in\left[\frac{1}{2c^{-}},1\right],\end{cases}$ (21)
such that
$h^{c}(p)=h\left(\min\left\{c^{-}p,c^{+}\left(1-p\right)\right\}\right).$ (22)
In conclusion,
$H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})=\mathbb{E}\left[h\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right]\leq h\left(\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]\right)\leq h\left(P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})\right),$ (23)
which implies
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=h^{-1}\left(H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)\leq P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f}).$ (24)